OpenStudy (anonymous):

the width of a rectangle is 7 feet less its length. If the perimeter of a rectangle is 58 feet, find the lenght & width of the rectangle

OpenStudy (anonymous):

W=L-7 2W+2L=58 2*(L-7)+2L=58 2L-14+2L=58 4L=72 L=18 in W=11 in